Control Chart Chocolate Viscosity 570 540 530 520 10 Subgroup # Control Chart-Chocolate Viscosity 50 s45 40 35 25 20 15 10 10 15 20 25 Subgroup #11. An operator working at a company that makes chocolate syrup rocorded the ollowing d on the syrup viscosity, using subgroups.of.10samples Answer the following questions 114 marks subgroupl average (cl) s (cP) 534 545 561 576 523 571 547 584 552 541 545 27 25 23 24 28 10 12 13 26 24 29 26 28 23 551 522 579 549 508 569 574 563 561 548 556 553 13798 15 16 17 19 20 21 28 23 23 24 25 sum 27 23 670 (A) Determine the central lines and control limits, and add these to the control charts below (label each line). Is the process in control? (B) The company management team has set internal specifications of 530+90 cP. As it is now is the process capable? (C) How does the Cpk compare to the Cp value? What does this indicate? (D) Product that is too viscous must be reworked (this costs SS). With the current process, how many units of product per 1000 fall above the USL?
Answer #1

Solution

Back-up Theory

Let Xdouble bar = (1/k)∑(i = 1, k)xibar, and sbar = (1/k)∑(i = 1, k)si where xibar = mean and si = standard deviation of the ith sub-group

The Xbar part of Xbar-s Chart has:

Central line: CL = Xdouble bar

Lower Control Limit: LCL = Xdouble bar – A3sbar

Upper Control Limit: UCL = Xdouble bar + A3sbar;

The s part of Xbar-s Chart has:

Central line: CL = sbar

Lower Control Limit: LCL = B3sbar

Upper Control Limit: UCL = B4sbar, where

A3, B3, B4 are constants which can be directly obtained from standard Control Chart Constants Table.

Now to work out the solution,

Only final results are given below. Detailed working follow at the end.

Part (A)

Final CL and Control Limits are:

Xbar Chart

CL

554.1

LCL

528.7988

UCL

579.4013

sChart

CL

25.95

LCL

7.3698

UCL

44.5302

ANSWER 1

At the first stage 5 sub-groups were found to be out of control limits. Reworking the limits after eliminating these sub-groups, saw all remaining sub-groups within limits.

Since only 20% of the sub-groups were out of limits, the process is considered ‘in control’. ANSWER 2

Part (B)

Given 530 ± 90, tolerance band is 180.

Estimate of process standard deviation = sbar = 25.95 and (6 x sbar) = 155.7.

Since (6 x sbar) < tolerance band, the process is considered ‘capable of meeting the specification’.

However, since process centering being at 554, which is much above the center of the specification band, despite being capable process, rejections are likely. ANSWER

Part (C)

Cpk compares well with the cp indicating in-built process capability. Nevertheless, the actual performance would depend on the process centering. ANSWER

Part (D)

With Xdoublebar at 554.1, sbar at 25.95 and USL at 620, probability of rework

= P[Z > {(620 – 554.1)/25.95}], where Z ~ N(0, 1)

= P(Z > 2.539)

= 0.0056.[using Excel Function: Statistical NORMSDIST]

So, number of units requiring rework in a lot of 1000 = 1000 x 0.0056

= 5.6 ~ 6 ANSWER

DONE

Details of working

k

25

n

10

A3

0.975

B3

0.284

B4

1.716

∑xibar

13798

∑si

670

Xdouble bar

551.92

sbar

26.8

Xbar Chart

CL

551.92

LCL

525.79

UCL

578.05

sChart

CL

26.8

LCL

7.6112

UCL

45.9888

Subgroups 6,9,15,16,

18 are out of control

limits.

Deleting these sub-

groups

k

20

n

10

A3

0.975

B3

0.284

B4

1.716

∑xibar

11082

∑si

519

Xdouble bar

554.1

sbar

25.95

Xbar Chart

CL

554.1

LCL

528.7988

UCL

579.4013

sChart

CL

25.95

LCL

7.3698

UCL

44.5302

Now all sub-groups are

within limits.